OpenStudy (anonymous):

find the vertical, horizontal and oblique asymptotes of x^2/(x=1)<-(under radical i'm just not sure how to type the radical in)

OpenStudy (anonymous):

So you have \[x^2 \over \sqrt{x-1}\]

OpenStudy (anonymous):

Yes! and i need the vertical, horizontal, and oblique asymptotes

OpenStudy (anonymous):

Vertical: Vertical asymptotes occur when the denominator is equal to zero. So \[\sqrt{x-1} = 0\rightarrow x = 1\] We have a vertical asymptote at x = 1. Horizontal: To find the location of the horizontal asymptote we have to look at the degree of the numerator and denominator. Let n be the degree of the numerator (n = 2) and m be the degree of the denominator (m= 1/2). Since n>m, there is no horizontal asymptote. If n=m then y location of the ratio of leading coefficients. (an/am) and if n<m there is a horizontal asymptote at y = 0. Oblique: Oblique asymptotes exist when n=m+1. This function has no oblique asymptotes. If an oblique asymptote exist, perform long division and discard the remainder to find the equation of the oblique asymptote.

OpenStudy (anonymous):

thank you!